# Towards Robust Influence Functions with Flat Validation Minima

### v2

@@ 1 Introduction
Training data is the fundamental component of machine learning, and its quality is critical for ensuring accurate and reliable predictions (Algan & Ulusoy, 2021; de Vargas et al., 2023). As large-scale foundation models, such as large language models (LLMs) (Brown et al., 2020) and text-to-image generative models (Rombach et al., 2022), become more prevalent, the demand for vast amounts of data to achieve high performance has grown exponentially. However, datasets—especially those sourced from the internet or generated by models—are often uncurated and may contain quite a few anomalous or mislabeled samples (Pleiss et al., 2020). These issues degrade model performance, making it crucial to understand how training data and trained models interact (Hammoudeh & Lowd, 2024). One key aspect of this interaction is assessing the influence of individual training samples on model predictions (Koh & Liang, 2017).
@@
A widely used approach for this purpose is the Influence Function (IF) (Koh & Liang, 2017), which provides insights into the predictions of Deep Neural Networks by estimating the impact of individual training samples. Specifically, IF estimates the change in the empirical risk on the validation set (or loss on a specific validation sample) resulting from small perturbations to the training set (e.g., the removal of a single training sample), using an estimated parameter change as a proxy. By serving as an efficient surrogate for the computationally expensive leave-one-out (LOO) retraining approach (Cook & Weisberg, 1982), IF enables the identification of the most influential samples in the training set of deep networks without retraining for each sample. Recent studies have successfully leveraged IF across various tasks, demonstrating its strong potential in practice  (Pruthi et al., 2020; Han et al., 2020; Kong et al., 2022; Kim et al., 2023; Kwon et al., 2024).
Despite its broad applicability, existing Influence Functions often struggle when applied to noisy training data, posing a significant challenge since IF is supposed to distinguish high-quality data from noisy or anomalous samples. Specifically, we observe that regardless of whether a first-order (Pruthi et al., 2020) or second-order (Koh & Liang, 2017) approximation is used to estimate parameter change, the estimated influence tends to be ineffective once the model has converged on the noisy training set, as illustrated in Figure 2. This suggests that the limitation does not originate from inaccuracies in parameter change estimation, such as errors in inverse Hessian approximation, which have been the primary focus of previous studies (Koh & Liang, 2017; Kwon et al., 2024). Instead, we provide a novel understanding that the issue lies in the estimation of loss change, particularly the impact of local sharpness in the validation risk. Figure 1 provides an intuitive visualization of how local sharpness affects Influence Function estimation. As shown, even with a perfectly estimated parameter change $\Delta\theta$, a sharp local region can still introduce a substantial gap between the estimated and actual influence, ultimately undermining the reliability of influence estimation.
@@ 2 Preliminaries
❶ The first step addresses the parameter change $\theta^{\star}_{z_{\text{tr}}}-\theta^{\star}$. Following Koh & Liang (2017), this parameter change can be approximated by a Newton step:
@@
❷ The second step estimates the change in the loss $\ell(z_{\text{val}},\theta^{\star}_{z_{\text{tr}}})-\ell(z_{\text{val}},\theta^{\star})$ of validation sample $z_{\text{val}}$ with respect to the parameter change $\theta^{\star}_{z_{\text{tr}}}-\theta^{\star}$. Using a first-order approximation, Koh & Liang (2017) approximate this change as:
@@ 3 Flat Validation Minima for Influence Estimation
Figure 2: Validation set accuracy and influence estimation performance (measured by ROC AUC) for the mislabeled samples detection task across training epochs. The orange and green curves represent the performance of the Influence Function in identifying mislabeled samples, which are assumed to have a negative influence on the clean validation set. Influence is estimated using LiSSA (Koh & Liang, 2017) for second-order parameter change approximation and TracIn (Pruthi et al., 2020) for first-order parameter change approximation. As shown, influence estimation performance is highly correlated with validation set accuracy, regardless of whether the parameter change is estimated using a first-order or second-order approximation. The experiment is conducted on the CIFAR-10N (Wei et al., 2022) dataset under the “worst” setting. For further details, see Section 4.1.
@@ 3.3 Flat Validation Minima via Sharpness-Aware Minimization
Specifically, this formulation naturally aligns with Sharpness-Aware Minimization (SAM) (Foret et al., 2021), a learning paradigm distinct from ERM. Instead of directly minimizing the empirical risk, SAM simultaneously minimizes both the loss value and the loss sharpness, aiming to find a flat minima. Intuitively, a flat minima leads to a model with better robustness and generalization ability.
@@
Figure 3: (a) The influence estimation performance, measured by ROC AUC, for the mislabeled sample detection task across tuning steps. The standard influence is estimated using LiSSA (Koh & Liang, 2017), with SAM (Foret et al., 2021) employed as the flat minima solver. As shown, the performance of the standard Influence Function decreases as $\hat{R}^{\gamma}_{\text{val}}(\theta)$ decreases, while our proposed Influence Function shows an improvement. The experiment is conducted on the CIFAR-10N (Wei et al., 2022) dataset under the “worst” setting. For more details, see Section 4.1. (b) Box plot of influence for clean samples estimated using the standard Influence Function across tuning steps. As illustrated, the absolute value of the estimated influence continuously decreases.
@@ 3.4 Influence Function in Flat Validation Minima
❷ For loss change, the standard IF utilizes a first-order approximation term for estimation. However, previous studies (Kim et al., 2023) have shown that the gradients of pre-trained neural networks can be approximated as a Gaussian distribution centered at zero. This property naturally causes the estimated influence, $\mathbb{E}_{z_{\text{tr}}\in S_{\text{tr}}}[\mathcal{I}(z_{\text{tr}},S_{\text{val}})]=\mathbb{E}_{z_{tr}\in S_{\text{tr}}}[g_{\text{val}}^{\top}H_{\text{tr}}^{-1}g_{z_{\text{tr}}}]$, to approach zero. Consequently, this leads to $\mu\to 0$, which severely limits the effectiveness of the standard estimation.
@@ 4.1 Mislabeled Sample Detection
We compare our proposed methods against several influence estimation approaches, including LiSSA (Koh & Liang, 2017), TracIn (Pruthi et al., 2020), GEX (Kim et al., 2023), and DataInf (Kwon et al., 2024). Additionally, we compare our methods with modified versions of LiSSA and DataInf, computed using the checkpoint from the training process that achieves the best results, referred to as LiSSA* and DataInf*. The experiments are conducted on CIFAR-10/-100N datasets (Wei et al., 2022). Detailed descriptions of the experimental setup can be found in Section B.1.
@@ 4.3 Influential Sample Identification in Text Generation
Table 4: ROC AUC (%) and Recall (%) for identifying influential samples in text generation tasks across different influence estimation methods. Bold indicates the best results and underline denotes the second-best results. The results for TracIn and DataInf were generated by executing the official implementation provided by DataInf. According to DataInf (Kwon et al., 2024), obtaining results around 99.99% is considered reasonable for this experiment.
@@
To further demonstrate the effectiveness of our proposed method, we evaluate its ability to identify influential samples in text generation, following Kwon et al. (2024). The assumption for this task is analogous to that in the mislabeled sample detection task (Section 4.1): for a generated sample, we assume that training samples from the same class contribute the most to the generation, while samples from other classes contribute less. Consequently, training samples that share the same class as the validation sample are expected to have a greater influence.
@@
We compare our proposed method against TracIn (Pruthi et al., 2020) and DataInf (Kwon et al., 2024). The experiment is conducted by fine-tuning LoRA (Hu et al., 2022) on the Llama-2-13B-chat model (Touvron et al., 2023) for the text-to-text generation task. We evaluate two tasks in text generation task: (i) Sentence transformations, and (ii) Math word problems (with reasoning). The detailed description for each task and dataset is given in Section B.2.
The results are presented in Table 4. Our proposed methods, VM and FVM consistently achieve the highest performance across both tasks. Specifically, we note that FVM does not always achieve the highest scores. This can be attributed to the fact that our fine-tuning process employed a simplified version of the flat minima solver, LPF (Bisla et al., 2022), using $M=1$ instead of the recommended $M=8$ in the original work. This suboptimal choice of hyperparameters may have limited the full potential of FVM, preventing it from consistently outperforming all baselines. Nonetheless, the overall trend still demonstrates the robustness and effectiveness of our method in identifying influential samples across text generation tasks.
@@ 4.4 Influential Sample Identification in Image Generation
Following Kwon et al. (2024), we further evaluate our proposed influence estimation method for identifying influential samples in image generation tasks, using the same experimental settings as described in Section B.2. We evaluate two tasks in text-to-image generation: (i) style generation and (ii) subject generation. For the style generation task, we combine three publicly available image-text pair datasets, each representing a different style: cartoons (Norod78, 2023), pixel-art (Jainr3, 2023), and line sketches (Zoheb, 2023). Further details of the experimental setup are provided in Section B.3.
@@ 4.5 Ablation Study
We begin with an ablation study to evaluate the effectiveness of the two key components of our proposed Influence Function: loss change and parameter change. For loss change, we compare the term $g_{z_{\text{val}}}^{\top}\Delta\theta$ used in the standard IF, referred to as “standard,” with $\Delta\theta^{\top}\tilde{H}_{\text{val}}\Delta\theta$ used in our proposed IF, referred to as “ours.” For parameter change, we compare $\epsilon H_{\text{tr}}^{-1}g_{z_{\text{tr}}}$ as “standard” with $\epsilon\tilde{H}^{-1}_{\text{val}}\tilde{g}_{z_{\text{tr}}}$ as “ours.” We conduct experiments on CIFAR-10N dataset under the “worst” setting. The model is tuned on the validation set with SAM (Foret et al., 2021) used as the flat minima solver, and the experimental setup is consistent with the details provided in Section 4.1.
@@ B.1 Mislabeled Sample Detection and Training Sample Relabeling
Datasets. Our evaluation is conducted on the CIFAR-10N and CIFAR-100N datasets (Wei et al., 2022), which are real-world noisy label variants of the CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009). CIFAR-10N includes five noisy label settings: “Aggregate”, “Random 1”, “Random 2”, “Random 3” and “Worst.” CIFAR-100N provides a single noisy label setting, “Noisy.” These noisy labels arise from human annotation errors. For our evaluation, we use the “Aggregate”, “Random 1” and “Worst” settings for CIFAR-10N, and the “Noisy” setting for CIFAR-100N.
Training. We follow the experiment settings in Wei et al. (2022). The ResNet-34 (He et al., 2016) is utilized for both CIFAR-10N and CIFAR-100N. The basic hyper-parameters settings are listed as follows: minibatch size (128), optimizer (SGD), initial learning rate (0.1), momentum (0.9), weight decay (0.0005), number of epochs (100), and learning rate decay (0.1 at 50 epochs). Standard data augmentation, including random crop and random horizontal flip, is applied to each dataset.
@@
For our proposed VM and FVM, to obtain $\tilde{\theta}$, we tune the trained model on the validation set with the following hyper-parameter settings: minibatch size (128), optimizer (SGD), initial learning rate (0.01), momentum (0.9), weight decay (0.0005), number of steps (1000), and learning rate decay (cosine). For FVM, SAM (Foret et al., 2021) is used as the flat minima solver, with the hyperparameter $\gamma$ set to 0.05 for CIFAR-10N and 0.1 for CIFAR-100N, in accordance with the original paper. In this experiment, we treat the samples with large influence scores as detected noisy samples.
@@ B.2 Text-To-Text Generation
Datasets. Following Kwon et al. (2024), we consider three text-to-text generation tasks: sentence transformations and math problems (w/ reasoning).
  • Sentence transformations: The model is tasked with applying a specific transformation to a given sentence. Further details can be found in Table 7.
  • Math problems (w/ reasoning): The model is presented with an arithmetic word problem and is expected to provide an intermediate reasoning step and a direct numerical answer before arriving at the final answer. A detailed description of the prompts and templates is provided in Table 8.
@@
Table 7: Description of the sentence transformations task prompts and templates. We consider 10 different types of sentence transformations, following the setup of Kwon et al. (2024).
Sentence transformations Example transformation of “Sunrises herald hopeful tomorrows”:
Reverse Order of Words tomorrows. hopeful herald Sunrises
Capitalize Every Other Letter sUnRiSeS hErAlD hOpEfUl tOmOrRoWs.
Insert Number 1 Between Every Word Sunrises 1herald 1hopeful 1tomorrows.
Replace Vowels with * S*nr*s*s h*r*ld h*p*f*l t*m*rr*ws.
Double Every Consonant SSunrriisseess hheralld hhopeffull ttomorrows.
Capitalize Every Word Sunrises Herald Hopeful Tomorrows.
Remove All Vowels Snrss hrld hpfl tmrrws.
Add ’ly’ To End of Each Word Sunrisesly heraldly hopefully tomorrows.ly
Remove All Consonants uie ea oeu ooo.
Repeat Each Word Twice Sunrises Sunrises herald herald hopeful hopeful tomorrows. tomorrows.
Table 8: Description of the math problem task templates. We consider 10 different types of math word problems, following the setup of Kwon et al. (2024).
Math word problems Template prompt question
Remaining pizza slices Lisa ate A slices of pizza and her brother ate B slices from a pizza that originally had C slices. How many slices of the pizza are left?
Reason: Combined slices eaten = A + B. Left = C - (A + B).
Chaperones needed for trip For every A students going on a field trip, there are B adults needed as chaperones. If C students are attending, how many adults are needed?
Reason: Adults needed = (B * C) // A.
Total number after purchase In an aquarium, there are A sharks and B dolphins. If they bought C more sharks, how many sharks would be there in total?
Reason: Total sharks = A + C.
Total game points Michael scored A points in the first game, B points in the second, C in the third, and D in the fourth game. What is his total points?
Reason: Total points = A + B + C + D.
Total reading hours Emily reads for A hours each day. How many hours does she read in total in B days?
Reason: Total hours read = A * B.
Shirt cost after discount A shirt costs A. There’s a B-dollar off sale. How much does the shirt cost after the discount?
Reason: Cost after discount = A - B.
Area of a garden A rectangular garden has a length of A meters and a width of B meters. What is its area?
Reason: Area = A * B.
Total savings If Jake saves A each week, how much will he save after B weeks?
Reason: Total savings = A * B.
Number of cupcake boxes A bakery sells cupcakes in boxes of A. If they have B cupcakes, how many boxes can they fill?
Reason: Boxes filled = B // A.
Interest earned John invests A at an annual interest rate of B%. How much interest will he earn after C years?
Reason: Interest = (A * B * C) // 100.
Training. We use LoRA (Hu et al., 2022) to Llama-2-13B-chat (Touvron et al., 2023). We apply LoRA to every query and value matrix of the attention layer in the Llama-2-13B-chat model. The basic hyper-parameters settings are listed as follows: minibatch size (64), optimizer (AdamW (Loshchilov & Hutter, 2019)), initial learning rate ($3\times 10^{-4}$), number of epochs (10), and learning rate decay (cosine). For LoRA hyperparameters, we set $r=8$ and $\alpha=32$. The training was performed using the HuggingFace Peft library (Mangrulkar et al., 2022)
@@
For TracIn, we use the final checkpoint for influence computation, following Kwon et al. (2024).
@@
For our proposed VM and FVM, to obtain $\tilde{\theta}$, we further fine-tune the previously tuned LoRA model (i.e., the model obtained after the initial fine-tuning on the training set) on the validation set for 5 epochs with an initial learning rate of $3\times 10^{-5}$. For FVM, we set $\epsilon=0.01$. LPF (Bisla et al., 2022) is employed as the flat minima solver, with the hyper-parameters $M=1$ and $\gamma=0.001$ for all tasks, in accordance with the original paper.
Table 9: Description of the text-to-image generation task templates. Each style has 200 training image-text pairs and 150 validation image-text pairs, following the setup of Kwon et al. (2024).
Image style Text prompt
Cartoon Generate an image in a specific cartoon style. A text sequence of the original dataset which describes an image.
Pixel Art Generate an image in a specific pixelart style. A text sequence of the original dataset which describes an image.
Sketch Generate an image in a specific black and white line sketch style. A text sequence of the original dataset which describes an image.

### v3

@@ 1 Introduction
Training data is the fundamental component of machine learning, and its quality is critical for ensuring accurate and reliable predictions (DBLP:journals/kbs/AlganU21:NoisyLabelsSurvey; DBLP:journals/kais/VargasACPB23:ImbalancedDataPreprocessing). As large-scale foundation models, such as large language models (LLMs) (DBLP:conf/nips/BrownMRSKDNSSAA20:LLM) and text-to-image generative models (DBLP:conf/cvpr/RombachBLEO22:StableDiffusion), become more prevalent, the demand for vast amounts of data to achieve high performance has grown exponentially. However, datasets—especially those sourced from the internet or generated by models—are often uncurated and may contain quite a few anomalous or mislabeled samples (DBLP:conf/nips/Pleiss0EW20:IdentifyingMislabeledData). These issues degrade model performance, making it crucial to understand how training data and trained models interact (DBLP:journals/ml/HammoudehL24:TrainingDataInfluenceSurvey). One key aspect of this interaction is assessing the influence of individual training samples on model predictions (DBLP:conf/icml/KohL17:IF).
@@
A widely used approach for this purpose is the Influence Function (IF) (DBLP:conf/icml/KohL17:IF), which provides insights into the predictions of Deep Neural Networks by estimating the impact of individual training samples. Specifically, IF estimates the change in the empirical risk on the validation set (or loss on a specific validation sample) resulting from small perturbations to the training set (e.g., the removal of a single training sample), using an estimated parameter change as a proxy. By serving as an efficient surrogate for the computationally expensive leave-one-out (LOO) retraining approach (cook1982ResidualsInfluenceInRegression), IF enables the identification of the most influential samples in the training set of deep networks without retraining for each sample. Recent studies have successfully leveraged IF across various tasks, demonstrating its strong potential in practice  (DBLP:conf/nips/PruthiLKS20:TraceIn; DBLP:conf/acl/HanWT20:ExplainingBlackBoxPredictions; DBLP:conf/iclr/KongSH22:IFonNoisyLabelLearning; DBLP:conf/nips/KimKY23:GEX; DBLP:conf/iclr/KwonWW024:DataInf).
Despite its broad applicability, existing Influence Functions often struggle when applied to noisy training data, posing a significant challenge since IF is supposed to distinguish high-quality data from noisy or anomalous samples. Specifically, we observe that regardless of whether a first-order (DBLP:conf/nips/PruthiLKS20:TraceIn) or second-order (DBLP:conf/icml/KohL17:IF) approximation is used to estimate parameter change, the estimated influence tends to be ineffective once the model has converged on the noisy training set, as illustrated in Figure 2. This suggests that the limitation does not originate from inaccuracies in parameter change estimation, such as errors in inverse Hessian approximation, which have been the primary focus of previous studies (DBLP:conf/icml/KohL17:IF; DBLP:conf/iclr/KwonWW024:DataInf). Instead, we provide a novel understanding that the issue lies in the estimation of loss change, particularly the impact of local sharpness in the validation risk. Figure 1 provides an intuitive visualization of how local sharpness affects Influence Function estimation. As shown, even with a perfectly estimated parameter change $\Delta\theta$, a sharp local region can still introduce a substantial gap between the estimated and actual influence, ultimately undermining the reliability of influence estimation.
@@ 2 Preliminaries
❶ The first step addresses the parameter change $\theta^{\star}_{z_{\text{tr}}}-\theta^{\star}$. Following DBLP:conf/icml/KohL17:IF, this parameter change can be approximated by a Newton step:
@@
❷ The second step estimates the change in the loss $\ell(z_{\text{val}},\theta^{\star}_{z_{\text{tr}}})-\ell(z_{\text{val}},\theta^{\star})$ of validation sample $z_{\text{val}}$ with respect to the parameter change $\theta^{\star}_{z_{\text{tr}}}-\theta^{\star}$. Using a first-order approximation, DBLP:conf/icml/KohL17:IF approximate this change as:
@@ 3 Flat Validation Minima for Influence Estimation
Figure 2: Validation set accuracy and influence estimation performance (measured by ROC AUC) for the mislabeled samples detection task across training epochs. The orange and green curves represent the performance of the Influence Function in identifying mislabeled samples, which are assumed to have a negative influence on the clean validation set. Influence is estimated using LiSSA (DBLP:conf/icml/KohL17:IF) for second-order parameter change approximation and TracIn (DBLP:conf/nips/PruthiLKS20:TraceIn) for first-order parameter change approximation. As shown, influence estimation performance is highly correlated with validation set accuracy, regardless of whether the parameter change is estimated using a first-order or second-order approximation. The experiment is conducted on the CIFAR-10N (DBLP:conf/iclr/WeiZ0L0022:CIFAR-10N/-100N) dataset under the “worst” setting. For further details, see Section 4.1.
@@ 3.3 Flat Validation Minima via Sharpness-Aware Minimization
Specifically, this formulation naturally aligns with Sharpness-Aware Minimization (SAM) (DBLP:conf/iclr/ForetKMN21:SAM), a learning paradigm distinct from ERM. Instead of directly minimizing the empirical risk, SAM simultaneously minimizes both the loss value and the loss sharpness, aiming to find a flat minima. Intuitively, a flat minima leads to a model with better robustness and generalization ability.
@@
Figure 3: (a) The influence estimation performance, measured by ROC AUC, for the mislabeled sample detection task across tuning steps. The standard influence is estimated using LiSSA (DBLP:conf/icml/KohL17:IF), with SAM (DBLP:conf/iclr/ForetKMN21:SAM) employed as the flat minima solver. As shown, the performance of the standard Influence Function decreases as $\hat{R}^{\gamma}_{\text{val}}(\theta)$ decreases, while our proposed Influence Function shows an improvement. The experiment is conducted on the CIFAR-10N (DBLP:conf/iclr/WeiZ0L0022:CIFAR-10N/-100N) dataset under the “worst” setting. For more details, see Section 4.1. (b) Box plot of influence for clean samples estimated using the standard Influence Function across tuning steps. As illustrated, the absolute value of the estimated influence continuously decreases.
@@ 3.4 Influence Function in Flat Validation Minima
❷ For loss change, the standard IF utilizes a first-order approximation term for estimation. However, previous studies (DBLP:conf/nips/KimKY23:GEX) have shown that the gradients of pre-trained neural networks can be approximated as a Gaussian distribution centered at zero. This property naturally causes the estimated influence, $\mathbb{E}_{z_{\text{tr}}\in S_{\text{tr}}}[\mathcal{I}(z_{\text{tr}},S_{\text{val}})]=\mathbb{E}_{z_{tr}\in S_{\text{tr}}}[g_{\text{val}}^{\top}H_{\text{tr}}^{-1}g_{z_{\text{tr}}}]$, to approach zero. Consequently, this leads to $\mu\to 0$, which severely limits the effectiveness of the standard estimation.
@@ 4.1 Mislabeled Sample Detection
We compare our proposed methods against several influence estimation approaches, including LiSSA (DBLP:conf/icml/KohL17:IF), TracIn (DBLP:conf/nips/PruthiLKS20:TraceIn), GEX (DBLP:conf/nips/KimKY23:GEX), and DataInf (DBLP:conf/iclr/KwonWW024:DataInf). Additionally, we compare our methods with modified versions of LiSSA and DataInf, computed using the checkpoint from the training process that achieves the best results, referred to as LiSSA* and DataInf*. The experiments are conducted on CIFAR-10/-100N datasets (DBLP:conf/iclr/WeiZ0L0022:CIFAR-10N/-100N). Detailed descriptions of the experimental setup can be found in Section B.1.
@@ 4.3 Influential Sample Identification in Text Generation
Table 4: ROC AUC (%) and Recall (%) for identifying influential samples in text generation tasks across different influence estimation methods. Bold indicates the best results and underline denotes the second-best results. The results for TracIn and DataInf were generated by executing the official implementation provided by DataInf. According to DataInf (DBLP:conf/iclr/KwonWW024:DataInf), obtaining results around 99.99% is considered reasonable for this experiment.
@@
To further demonstrate the effectiveness of our proposed method, we evaluate its ability to identify influential samples in text generation, following DBLP:conf/iclr/KwonWW024:DataInf. The assumption for this task is analogous to that in the mislabeled sample detection task (Section 4.1): for a generated sample, we assume that training samples from the same class contribute the most to the generation, while samples from other classes contribute less. Consequently, training samples that share the same class as the validation sample are expected to have a greater influence.
@@
We compare our proposed method against TracIn (DBLP:conf/nips/PruthiLKS20:TraceIn) and DataInf (DBLP:conf/iclr/KwonWW024:DataInf). The experiment is conducted by fine-tuning LoRA (DBLP:conf/iclr/HuSWALWWC22:LoRA) on the Llama-2-13B-chat model (touvron2023llama2) for the text-to-text generation task. We evaluate two tasks in text generation task: (i) Sentence transformations, and (ii) Math word problems (with reasoning). The detailed description for each task and dataset is given in Section B.2.
The results are presented in Table 4. Our proposed methods, VM and FVM consistently achieve the highest performance across both tasks. Specifically, we note that FVM does not always achieve the highest scores. This can be attributed to the fact that our fine-tuning process employed a simplified version of the flat minima solver, LPF (DBLP:conf/aistats/BislaWC22:LPF-SGD), using $M=1$ instead of the recommended $M=8$ in the original work. This suboptimal choice of hyperparameters may have limited the full potential of FVM, preventing it from consistently outperforming all baselines. Nonetheless, the overall trend still demonstrates the robustness and effectiveness of our method in identifying influential samples across text generation tasks.
@@ 4.4 Influential Sample Identification in Image Generation
Following DBLP:conf/iclr/KwonWW024:DataInf, we further evaluate our proposed influence estimation method for identifying influential samples in image generation tasks, using the same experimental settings as described in Section B.2. We evaluate two tasks in text-to-image generation: (i) style generation and (ii) subject generation. For the style generation task, we combine three publicly available image-text pair datasets, each representing a different style: cartoons (huggingfaceNorod78cartoonblipcaptionsDatasets), pixel-art (huggingfaceJainr3diffusiondbpixelartDatasets), and line sketches (huggingfaceZohebsketchsceneDatasets). Further details of the experimental setup are provided in Section B.3.
@@ 4.5 Ablation Study
We begin with an ablation study to evaluate the effectiveness of the two key components of our proposed Influence Function: loss change and parameter change. For loss change, we compare the term $g_{z_{\text{val}}}^{\top}\Delta\theta$ used in the standard IF, referred to as “standard,” with $\Delta\theta^{\top}\tilde{H}_{\text{val}}\Delta\theta$ used in our proposed IF, referred to as “ours.” For parameter change, we compare $\epsilon H_{\text{tr}}^{-1}g_{z_{\text{tr}}}$ as “standard” with $\epsilon\tilde{H}^{-1}_{\text{val}}\tilde{g}_{z_{\text{tr}}}$ as “ours.” We conduct experiments on CIFAR-10N dataset under the “worst” setting. The model is tuned on the validation set with SAM (DBLP:conf/iclr/ForetKMN21:SAM) used as the flat minima solver, and the experimental setup is consistent with the details provided in Section 4.1.
@@ B.1 Mislabeled Sample Detection and Training Sample Relabeling
Datasets. Our evaluation is conducted on the CIFAR-10N and CIFAR-100N datasets (DBLP:conf/iclr/WeiZ0L0022:CIFAR-10N/-100N), which are real-world noisy label variants of the CIFAR-10 and CIFAR-100 datasets (krizhevsky2009learning:CIFAR-10/-100). CIFAR-10N includes five noisy label settings: “Aggregate”, “Random 1”, “Random 2”, “Random 3” and “Worst.” CIFAR-100N provides a single noisy label setting, “Noisy.” These noisy labels arise from human annotation errors. For our evaluation, we use the “Aggregate”, “Random 1” and “Worst” settings for CIFAR-10N, and the “Noisy” setting for CIFAR-100N.
Training. We follow the experiment settings in DBLP:conf/iclr/WeiZ0L0022:CIFAR-10N/-100N. The ResNet-34 (DBLP:conf/cvpr/HeZRS16:ResNet) is utilized for both CIFAR-10N and CIFAR-100N. The basic hyper-parameters settings are listed as follows: minibatch size (128), optimizer (SGD), initial learning rate (0.1), momentum (0.9), weight decay (0.0005), number of epochs (100), and learning rate decay (0.1 at 50 epochs). Standard data augmentation, including random crop and random horizontal flip, is applied to each dataset.
@@
For our proposed VM and FVM, to obtain $\tilde{\theta}$, we tune the trained model on the validation set with the following hyper-parameter settings: minibatch size (128), optimizer (SGD), initial learning rate (0.01), momentum (0.9), weight decay (0.0005), number of steps (1000), and learning rate decay (cosine). For FVM, SAM (DBLP:conf/iclr/ForetKMN21:SAM) is used as the flat minima solver, with the hyperparameter $\gamma$ set to 0.05 for CIFAR-10N and 0.1 for CIFAR-100N, in accordance with the original paper. In this experiment, we treat the samples with large influence scores as detected noisy samples.
@@ B.2 Text-To-Text Generation
Datasets. Following DBLP:conf/iclr/KwonWW024:DataInf, we consider three text-to-text generation tasks: sentence transformations and math problems (w/ reasoning).
  • Sentence transformations: The model is tasked with applying a specific transformation to a given sentence. Further details can be found in DBLP:conf/iclr/KwonWW024:DataInf.
  • Math problems (w/ reasoning): The model is presented with an arithmetic word problem and is expected to provide an intermediate reasoning step and a direct numerical answer before arriving at the final answer. A detailed description of the prompts and templates can be found in DBLP:conf/iclr/KwonWW024:DataInf.
@@
Training. We use LoRA (DBLP:conf/iclr/HuSWALWWC22:LoRA) to Llama-2-13B-chat (touvron2023llama2). We apply LoRA to every query and value matrix of the attention layer in the Llama-2-13B-chat model. The basic hyper-parameters settings are listed as follows: minibatch size (64), optimizer (AdamW (DBLP:conf/iclr/LoshchilovH19:AdamW)), initial learning rate ($3\times 10^{-4}$), number of epochs (10), and learning rate decay (cosine). For LoRA hyperparameters, we set $r=8$ and $\alpha=32$. The training was performed using the HuggingFace Peft library (peft)
@@
For TracIn, we use the final checkpoint for influence computation, following DBLP:conf/iclr/KwonWW024:DataInf.
@@
For our proposed VM and FVM, to obtain $\tilde{\theta}$, we further fine-tune the previously tuned LoRA model (i.e., the model obtained after the initial fine-tuning on the training set) on the validation set for 5 epochs with an initial learning rate of $3\times 10^{-5}$. For FVM, we set $\epsilon=0.01$. LPF (DBLP:conf/aistats/BislaWC22:LPF-SGD) is employed as the flat minima solver, with the hyper-parameters $M=1$ and $\gamma=0.001$ for all tasks, in accordance with the original paper.